# Computing Hamiltonian Paths with Partial Order Restrictions

### v1

@@ 1 Introduction
In this paper, we will study the problem of finding (minimum) Hamiltonian paths with precedence controls. The focus on the structure of the edge set (as opposed to TSP) merits the following definition.
@@
Is there a Hamiltonian path $(v_{1},\dots,v_{n})$ in $G$ such that for all $i,j\in\{1,\dots,n\}$ it holds that if $(v_{i},v_{j})\in\pi$, then $i\leq j$?
If, in addition, we are given a cost function $c:E\rightarrow\mathbb{Q}$ the Minimum Partially Ordered Hamiltonian Path Problem (MinPOHPP) asks for an ordered Hamiltonian path $\mathcal{P}$ with minimum cost such that $\lambda(\mathcal{P})$ is a linear extension of $\pi$.
@@ 1.2 Our Contribution
The (Min)POHPP as well as the TSP-PC can be solved in time $\mathcal{O}(k^{2}n^{k})$ for arbitrary graphs with $n$ vertices if we restrict ourselves to partial orders of width $k$ [16]. We complement this result by showing that the (Min)POHPP and the TSP-PC are W[1]-hard if they are parameterized by the width of the partial order. Furthermore, we show that it is unlikely that the running time of the algorithm given in [16] can be improved significantly since there is no $f(k)n^{o(k)}$-time algorithm for (Min)POHPP or TSP-PC for any computable function $f$ assuming that the Exponential Time Hypothesis is true. Lastly, we show that on the class of outerplanar graphs the (Min)POHPP can be solved in time $\mathcal{O}(n^{2})$ for arbitrary partial orders.
@@ 2.2 Graphs
A graph is a split graph if its vertex set can be partitioned into a clique and an independent set. A graph is a complete split graph if there is such a partition where every vertex of the independent set is adjacent to all the vertices of the clique. A graph $G$ is a bipartite graph if its vertex set can be partitioned into two independent sets $A$ and $B$. Furthermore, a bipartite graph is complete bipartite if every vertex of $A$ is connected to every vertex of $B$. A bipartite graph is balanced if $|A|=|B|$.
@@
A graph is called planar if it has a crossing-free embedding in the plane, and together with this embedding it is called plane. For a plane graph $G$ we call the regions of $\mathbb{R}^{2}\setminus G$ the faces of $G$. Every plane graph has exactly one unbounded face which is called the outer face. A graph is called outerplanar if it has a crossing-free embedding such that all of the vertices belong to the outer face and such an embedding is also called outerplanar.
For a connected graph $G$, a vertex $v\in V(G)$ is a cut vertex of $G$ if $G-v$ is not connected. If a graph has no cut vertex, then it is called 2-connected. The blocks of a graph are its inclusion maximal 2-connected subgraphs. The block-cut tree $\cal T$ of $G$ is the bipartite graph that contains a vertex for every cut vertex of $G$ and a vertex for every block of $G$ and the vertex of block $B$ is adjacent to the vertex of a cut vertex $v$ in $\cal T$ if $B$ contains $v$. It is easy to see that the block-cut tree of a connected graph is in fact a tree.
3 $\mathsf{NP}$-completeness of POHPP
@@
Consider those complete split graphs with independent set $I$ and clique $C$ for which $|I|-1=|C|$. Obviously, any Hamiltonian path must alternate between $I$ and $C$. Thus, finding a Hamiltonian path in such a graph is equivalent to finding one in a complete bipartite graph, as none of the edges between vertices in $C$ can be used. It can easily be seen that it is not an issue that $|I|$ has one element more than $|C|$. Therefore, POHPP is $\mathsf{NP}$-complete on complete split graphs as well.
@@
The last two results imply that the POHPP is also $\mathsf{NP}$-hard for some other graph classes for which the regular Hamiltonian path problem can be solved in polynomial time. For example, it is known that the Hamiltonian path problem can even be solved dynamically in polynomial time on chain graphs, as well as on threshold graphs [10]. However, as chain graphs contain the complete bipartite graphs and threshold graphs contain the complete split graphs, the POHPP is $\mathsf{NP}$-complete on both of these classes. The same holds for the class of interval graphs for which Hamiltonian path can be solved in linear time [39] and which contain the threshold graphs.
@@
Partial Search Order Problem (PSOP) of graph search $\cal A$
@@
A graph $G$, a partial order on $V(G)$.
@@
Is there a search ordering of $G$ constructed by $\cal A$ that is a linear extension of $\pi$.
@@
We can also extend this result to the special variant LDFS which was introduced by Corneil and Krueger [19]. On complete bipartite graphs every DFS ordering is also an LDFS ordering (see [41] and [54, Theorem 1.2.74]). Therefore, Theorem 3.11 implies that the PSOP is also $\mathsf{NP}$-hard for LDFS on this graph class.
@@ 4 Partial Orders of Bounded Width
As we have seen in the last section, the POHPP is $\mathsf{NP}$-complete for partial orders even if the height of the partial order is 2. Here, we will consider partial orders of fixed width. We will see that, unless $\mathsf{P}=\mathsf{NP}$, the (Min)POHPP is not $\mathsf{NP}$-complete for any fixed width. If a partial order has width 1, then it is a linear ordering. In this case, the POHPP simply asks whether a given linear vertex ordering induces a Hamiltonian path in the graph. This problem can be solved straightforwardly in time $\mathcal{O}(n+m)$ on a graph with $n$ vertices and $m$ edges. In 1985, Colbourn and Pulleyblank [16] formulated an idea for an algorithm that solves the Minimum Setup Scheduling Problem for partial orders of width $k$ in time $\mathcal{O}(k^{2}n^{k})$. The Minimum Setup Scheduling Problem is in essence the same as a MinPOHPP for complete graphs (or the path variant of TSP-PC) showing that MinPOHPP is in XP for the width as parameter. So far, it has remained open whether this algorithm can be improved to an FPT algorithm. We show that this is unlikely by proving that the (Min)POHPP is W[1]-hard if it is parameterized by the width of the partial order. Our proof yields an even stronger bound on the running time, as it shows that there is no $f(k)n^{o(k)}$-time algorithm for (Min)POHPP assuming that the Exponential Time Hypothesis is true.
Before we prove the W[1]-hardness result, we present a complete pseudo-code as well as a thorough analysis both of the correctness and the running time of the $\mathcal{O}(k^{2}n^{k})$ algorithm since Colbourn and Pulleyblank [16] only give a brief sketch of the ideas underlying the algorithm.
@@ 17                   $v\leftarrow C_{\omega}[x_{\omega}]$;
We prove this claim inductively. First we consider the tuples $A=(x_{1},\ldots,x_{k},\omega)$ with $\sum_{i=1}^{k}x_{i}=1$, i.e., the respective ordered path contains exactly one element. The $M$-value of such a tuple is set to 0 if and only if the first element of $C_{\omega}$ is a minimal element of $\pi$, otherwise it is set to $\infty$ (see line 11 of Algorithm 1). Therefore, an ordered path $\mathcal{P}$ fulfilling the Conditions iiii exists if and only if $M[A]<\infty$. If this is the case, then the minimal cost of such a path is $0=M[A]$.
@@ 4.2 W[1]-hardness
While Theorem 4.1 shows that Algorithm 1 has polynomial running time for a fixed $k$, the factor $n^{k}$ prevents this from being an FPT algorithm. In the following, we will prove that the POHPP parameterized by the width of the poset is $\mathsf{W}[1]$-hard, by reducing the Multicolor Clique Problem to the POHPP. Using this reduction, we show that no algorithm exists which has a significantly better running time than Algorithm 1 unless the Exponential Time Hypothesis (ETH) fails. The ETH, formulated by Impagliazzo et al. [37], states that the 3-SAT problem cannot be solved in time $2^{o(n)}\cdot(n+m)^{k}$ for formulas with $n$ variables and $m$ clauses and any fixed integer $k$.
@@
Multicolor Clique Problem (MCP)
@@
The MCP was shown to be $\mathsf{W}[1]$-hard by Fellows et al. [27]. In fact, in [21, 43] the authors show the following result.
@@
Assuming the Exponential Time Hypothesis, there is no $f(k)n^{o(k)}$ time algorithm for the Multicolor Clique Problem for any computable function $f$.
For an instance $G$ of the MCP we can assume that all color classes are of the same size $q$ (as adding isolated vertices does not change the existence or otherwise of a multicolor clique) and we denote these vertices by $v^{i}_{1},\ldots,v^{i}_{q}$ for each color class $i\in\{1,\ldots,k\}$.
@@
The main theorem of this section is a direct consequence of Lemmas 4.6, 4.10 and 4.16, Theorem 4.5 as well as the W[1]-hardness of the Multicolor Clique Problem [27].
@@ 5 Outerplanar Graphs
We consider the block-cut tree ${\cal T}$ of $G$. It is easy to see that ${\cal T}$ is a path if $G$ has a Hamiltonian path. Let $(B_{1},\ldots,B_{k})$ be this path where the $B_{i}$ are the blocks of $G$. For every $B_{i}$ let $\pi_{i}$ be the restriction of $\pi$ to $B_{i}$. For every $i\in\{2,...,k-1\}$, we define the partial order $\pi_{i}^{+}$ and $\pi_{i}^{-}$ as follows. In $\pi_{i}^{+}$ we add all the tuples that force the cut vertex in $B_{i-1}\cap B_{i}$ to be the first vertex and the cut vertex in $B_{i-1}\cap B_{i}$ to be the last. In $\pi_{i}^{-}$ we do this the other way round. The partial order $\pi_{1}^{+}$ forces the cut vertex in $B_{1}$ to be the last vertex and $\pi_{1}^{-}$ forces it to be the first vertex. Similar, $\pi_{k}^{+}$ forces the cut vertex of $B_{k}$ to be the first vertex and $\pi_{k}^{-}$ forces it to be the last vertex. Now we solve the MinPOHPP for all blocks $B_{i}$, once for $\pi_{i}^{+}$ and once for $\pi_{i}^{-}$. It is easy to see that the minimum cost ordered Hamiltonian path $\mathcal{P}$ of $G$ fulfilling the constraints of $\pi$ consists of minimum cost ordered Hamiltonian paths either for each instance $(B_{i},\pi_{i}^{+})$ or for each instance $(B_{i},\pi_{i}^{-})$. Thus, solving these instances is enough to solve the problem for the whole graph $G$. Note that the blocks of $G$ are also elements of $\mathcal{G}$ since $\mathcal{G}$ is hereditary.

### v2

@@ 1 Introduction
In this paper, we will study the problem of finding (minimum) Hamiltonian paths with precedence constraints. As a Hamiltonian path implies two different linear orderings, we fix one of these orderings and call it an ordered Hamiltonian path. The focus on the structure of the edge set (as opposed to TSP) merits the following definition.
@@
Is there an ordered Hamiltonian path $(v_{1},\dots,v_{n})$ in $G$ such that for all $i,j\in\{1,\dots,n\}$ it holds that if $(v_{i},v_{j})\in\pi$, then $i\leq j$?
If, in addition, we are given a cost function $c:E\rightarrow\mathbb{Q}$ the Minimum Partially Ordered Hamiltonian Path Problem (MinPOHPP) asks for an ordered Hamiltonian path $\mathcal{P}$ with minimum cost such that the linear ordering of $\mathcal{P}$ is a linear extension of $\pi$.
@@ 1.2 Our Contribution
The (Min)POHPP as well as the TSP-PC can be solved in time $\mathcal{O}(k^{2}n^{k})$ for arbitrary graphs with $n$ vertices if we restrict ourselves to partial orders of width $k$ [16]. We complement this result by showing that the (Min)POHPP and the TSP-PC are $\mathsf{W}[1]$-hard if they are parameterized by the width of the partial order. Furthermore, we show that it is unlikely that the running time of the algorithm given in [16] can be improved significantly since there is no $f(k)n^{o(k)}$-time algorithm for (Min)POHPP or TSP-PC for any computable function $f$ assuming that the Exponential Time Hypothesis is true. In contrast, we present an $\mathsf{FPT}$ algorithm for another parameter, the distance to linear order. This parameter measures the number of vertices that have to be deleted from $G$ to make the partial order a linear order. Lastly, we show that on the class of outerplanar graphs the (Min)POHPP can be solved in time $\mathcal{O}(n^{2})$ for arbitrary partial orders.
@@ 2.2 Graphs
A graph is a split graph if its vertex set can be partitioned into a clique and an independent set. A graph is a complete split graph if there is such a partition where every vertex of the independent set is adjacent to all the vertices of the clique. A graph $G$ is a bipartite graph if its vertex set can be partitioned into two independent sets $A$ and $B$. Furthermore, a bipartite graph is complete bipartite if every vertex of $A$ is adjacent to every vertex of $B$. A bipartite graph is balanced if $|A|=|B|$.
@@
A graph is called planar if it has a crossing-free embedding in the plane, and together with this embedding it is called a plane graph. For a plane graph $G$ we call the regions of $\mathbb{R}^{2}\setminus G$ the faces of $G$. Every plane graph has exactly one unbounded face which is called the outer face. A graph is called outerplanar if it has a crossing-free embedding such that all of the vertices belong to the outer face and such an embedding is also called outerplanar.
For a connected graph $G$, a vertex $v\in V(G)$ is a cut vertex of $G$ if $G-v$ is not connected. If a graph has no cut vertex, then it is called 2-connected. The blocks of a graph are its inclusion maximal 2-connected subgraphs. The block-cut tree $\mathcal{T}$ of $G$ is the bipartite graph that contains a vertex for every cut vertex of $G$ and a vertex for every block of $G$ and the vertex of block $B$ is adjacent to the vertex of a cut vertex $v$ in $\mathcal{T}$ if $B$ contains $v$. It is easy to see that the block-cut tree of a connected graph is in fact a tree.
3 NP-completeness of POHPP
@@
Note that the POHPP stays $\mathsf{NP}$-complete even if one set of the bipartition of the complete bipartite graph has one vertex more than the other. As we have seen in the proof of Theorem 3.7, we can always assume that a proper Hamiltonian path starts in $A$ and ends in $B$. If we add one vertex to $A$ and force it to be to the right of all other vertices, then there is solution for the POHPP of the new instance if and only if there is such a solution for the old instance. We can use this observation to show that the POHPP is also $\mathsf{NP}$-complete on complete split graphs. Consider those complete split graphs with independent set $I$ and clique $C$ for which $|I|-1=|C|$. Obviously, any Hamiltonian path must alternate between $I$ and $C$. Thus, finding a Hamiltonian path in such a graph is equivalent to finding one in a complete bipartite graph, as none of the edges between vertices in $C$ can be used. Therefore, POHPP is $\mathsf{NP}$-complete on complete split graphs as well.
@@
The last two results imply that the POHPP is also $\mathsf{NP}$-hard for some other graph classes for which the regular Hamiltonian path problem can be solved in polynomial time. For example, it is known that the Hamiltonian path problem can even be solved dynamically in polynomial time on chain graphs, as well as on threshold graphs [10]. Here, the dynamic algorithm gives an update on the existence of a Hamiltonian path in constant time after addition or deletion of an edge. However, as chain graphs contain the complete bipartite graphs and threshold graphs contain the complete split graphs, the POHPP is $\mathsf{NP}$-complete on both of these classes. The same holds for the class of interval graphs for which Hamiltonian path can be solved in linear time [40] and which contain the threshold graphs.
@@
Partial Search Order Problem (PSOP) of graph search $\mathcal{A}$
@@
A graph $G$, a partial order $\pi$ on $V(G)$.
@@
Is there a search ordering of $G$ constructed by $\mathcal{A}$ that is a linear extension of $\pi$.
@@
We can also extend this result to the special variant LDFS which was introduced by Corneil and Krueger [19]. On complete bipartite graphs every DFS ordering is also an LDFS ordering (see [42] and [56, Theorem 1.2.74]). Therefore, Theorem 3.11 implies that the PSOP is also $\mathsf{NP}$-hard for LDFS on this graph class.
@@ 4 Partial Orders of Bounded Width
As we have seen in the last section, the POHPP is $\mathsf{NP}$-complete for partial orders even if the height of the partial order is 2. Here, we will consider partial orders of fixed width. We will see that, unless $\mathsf{P}=\mathsf{NP}$, the (Min)POHPP is not $\mathsf{NP}$-complete for any fixed width. If a partial order has width 1, then it is a linear ordering. In this case, the POHPP simply asks whether a given linear vertex ordering induces a Hamiltonian path in the graph. This problem can be solved straightforwardly in time $\mathcal{O}(n+m)$ on a graph with $n$ vertices and $m$ edges. In 1985, Colbourn and Pulleyblank [16] formulated an idea for an algorithm that solves the Minimum Setup Scheduling Problem for partial orders of width $k$ in time $\mathcal{O}(k^{2}n^{k})$. The Minimum Setup Scheduling Problem is in essence the same as a MinPOHPP for complete graphs (or the path variant of TSP-PC) showing that MinPOHPP is in $\mathsf{XP}$ for the width as parameter. So far, it has remained open whether this algorithm can be improved to an $\mathsf{FPT}$ algorithm. We show that this is unlikely by proving that the (Min)POHPP is $\mathsf{W}$[1]-hard if it is parameterized by the width of the partial order. Our proof yields an even stronger bound on the running time, as it shows that there is no $f(k)n^{o(k)}$-time algorithm for (Min)POHPP assuming that the Exponential Time Hypothesis is true.
Before we prove the $\mathsf{W}$[1]-hardness result, we present a complete pseudo-code as well as a thorough analysis both of the correctness and the running time of the $\mathcal{O}(k^{2}n^{k})$ algorithm since Colbourn and Pulleyblank [16] only give a brief sketch of the ideas underlying the algorithm.
@@ 17                   $v\leftarrow C_{\omega}[x_{\omega}]$;
We prove this claim inductively. First we consider the tuples $A=(x_{1},\ldots,x_{k},\omega)$ with $\sum_{i=1}^{k}x_{i}=1$, i.e., the respective ordered path contains exactly one element. The $M$-value of such a tuple is set to 0 if and only if the first element of $C_{\omega}$ is a minimal element of $\pi$, otherwise it is set to $\infty$ (see lines 11 of Algorithm 1). Therefore, an ordered path $\mathcal{P}$ fulfilling the Conditions iiii exists if and only if $M[A]<\infty$. If this is the case, then the minimal cost of such a path is $0=M[A]$.
@@ 4.2 W[1]-hardness
While Theorem 4.1 shows that Algorithm 1 has polynomial running time for a fixed $k$, the factor $n^{k}$ prevents this from being an $\mathsf{FPT}$ algorithm. In the following, we will prove that the POHPP parameterized by the width of the poset is $\mathsf{W}[1]$-hard, by reducing the Multicolored Clique Problem to the POHPP. Using this reduction, we show that no algorithm exists which has a significantly better running time than Algorithm 1 unless the Exponential Time Hypothesis (ETH) fails. The ETH, formulated by Impagliazzo et al. [38], states that the 3-SAT problem cannot be solved in time $2^{o(n)}\cdot(n+m)^{k}$ for formulas with $n$ variables and $m$ clauses and any fixed integer $k$.
@@
Multicolored Clique Problem (MCP)
@@
The MCP was shown to be $\mathsf{W}[1]$-hard by Pietrzak [51] and independently by Fellows et al. [27]. In fact, in [21, 44] the authors show the following result.
@@
Assuming the Exponential Time Hypothesis, there is no $f(k)n^{o(k)}$ time algorithm for the Multicolored Clique Problem for any computable function $f$.
For an instance $G$ of the MCP we can assume that all color classes are of the same size $q$ (as adding isolated vertices does not change the existence or otherwise of a multicolored clique) and we denote these vertices by $v^{i}_{1},\ldots,v^{i}_{q}$ for each color class $i\in\{1,\ldots,k\}$.
@@
The main theorem of this section is a direct consequence of Lemmas 4.6, 4.10 and 4.16, Theorem 4.5 as well as the $\mathsf{W}$[1]-hardness of the Multicolored Clique Problem [27].
@@
4.3 Another Parameter: Distance to Linear Order
As shown above, it is unlikely that we can give an $\mathsf{FPT}$-algorithm for the POHPP if it is parameterized by the width of the partial order. However, we can ask whether this bound on the width can somehow be meaningfully strengthened to achieve an $\mathsf{FPT}$-algorithm.
For a given partial order $\pi$ on a set $X$, we say the distance of $\pi$ to linear order is equal to $|X|-h(\pi)$ where $h(\pi)$ is the height of $\pi$. It is easy to see that a partial order with a distance to linear order of $k$ has width at most $k+1$. Contrary, partial orders of width 1 have distance to linear order 00. However, even partial orders of width $2$ could have a distance to linear order of $\frac{|X|}{2}$.
Theorem 4.20.
Given a graph $G$ with $n$ vertices and a partial order $\pi$ on $V(G)$ of distance to linear order $k\geq 1$, we can solve the MinPOHPP in time $\mathcal{O}(k^{3}2^{k}n+m+|\pi|)$.
Proof 4.21.
We first compute a maximum chain $C$ of $\pi$ which can be done in time $\mathcal{O}(|\pi|+n)$ [33]. Similarly as in Algorithm 1, we write $C[i]$ to describe the $i$-th element of the chain $C$, i.e., $C[i]\prec_{\pi}C[j]$ if and only if $i<j$. Let $S$ be the set of vertices of $G$ that are not contained in $C$. It holds that $|S|=k$. For every vertex $v\in V(G)$, we construct a list $A(v)$ that contains all neighbors of $v$ in $S$ together with the weight of the edge between $v$ and the neighbor. If $v=C[i]$, then $A(v)$ also contains $C[i-1]$ and the corresponding edge weight if $C[i-1]$ is a neighbor of $C[i]$ in $G$. Furthermore, we have a list $\Pi_{\prec}(v)$ that contains all $x\in S$ with $x\prec_{\pi}v$. For every element $v\in S$, we define the variable $\xi(v)$ that contains the maximal index $j$ for which $C[j]\prec_{\pi}v$, i.e., the first $j$ elements of the chain $C$ are smaller than $v$ in $\pi$ and all other vertices of $C$ are not smaller than $v$ in $\pi$. Note that we can compute these lists and these variables in time $\mathcal{O}(|\pi|)$.
Now, the algorithm works as follows. We have a vector $M$ that has an entry for every tuple $(Z,i,u)$ with $Z\subseteq S$, $i\in\{0,\dots,|C|\}$ and $u\in Z\cup\{C[i]\}$. This entry is the minimum cost of an ordered path $\mathcal{P}$ of $G$ fulfilling the following properties (or $\infty$ if no such path exists):
  1. (i)
    $\mathcal{P}$ contains the $j$-th element of $C$ if and only if $j\leq i$,
  2. (ii)
    $\mathcal{P}$ contains a vertex $v\in S$ if and only if $v\in Z$,
  3. (iii)
    $\lambda(\mathcal{P})$ is a prefix of a linear extension of $\pi$,
  4. (iv)
    the last element of $\lambda(\mathcal{P})$ is the $u$.
First, we set all entries to $\infty$. We fill in the entries of $M$ inductively via the value $i+|Z|$. For $i+|Z|=1$, we have two cases. The $M$-value of entry $(\emptyset,1,C[1])$ is set to 00 if and only if $\Pi_{\prec}(C[1])=\emptyset$. The $M$-values of entries ($(\{u\},0,u)$ with $u\in S$ are set to 00 if and only if $\Pi_{\prec}(C[1])=\emptyset$ and $\xi(u)=0$.
Now assume that we have computed all $M$-values of tuples $(Z,i,u)$ with $i+|Z|=\ell$. Let $(Z,i,u)$ be an entry with $i+|Z|=\ell+1$. If $u\in Z$, then we check for all entries $(Z\setminus\{u\},i,v)$ whether $u\in A(v)$, whether $\xi(u)\leq i$ and whether $\Pi_{\prec}(u)\setminus Z=\emptyset$. If $u=C[i]$, then we check for all entries $(Z,i-1,v)$ whether $\Pi_{\prec}(u)\setminus Z=\emptyset$ and whether $v\in A(u)$. We omit the proof that the algorithm works correctly as it follows along the same lines as the proof of Theorem 4.1.
For the running time, we first observe that for any entry of $M$, we have to check at most $k+1$ entries of $M$ and for any of those entries we can do these checks in time $\mathcal{O}(k)$ since $|A(v)|\leq k+1$ and $|\Pi_{\prec}(v)|\leq k$. Thus, the algorithm needs time $\mathcal{O}(k^{2})$ per entry of $M$. Since $M$ has $\mathcal{O}(k\cdot 2^{k}\cdot n)$ entries, the overall running time of the procedure is $\mathcal{O}(k^{3}\cdot 2^{k}\cdot n)$. Combining this with the costs of computing $C$, $A$, and $\Pi_{\prec}$, we get the running time $\mathcal{O}(k^{3}\cdot 2^{k}\cdot n+m+|\pi|)$.
@@ 5 Outerplanar Graphs
We consider the block-cut tree ${\cal T}$ of $G$. It is easy to see that ${\cal T}$ is a path if $G$ has a Hamiltonian path. Let $(B_{1},\ldots,B_{k})$ be this path where the $B_{i}$ are the blocks of $G$. For every $B_{i}$ let $\pi_{i}$ be the restriction of $\pi$ to $B_{i}$. For every $i\in\{2,\dots,k-1\}$, we define the partial order $\pi_{i}^{+}$ and $\pi_{i}^{-}$ as follows. In $\pi_{i}^{+}$ we add all the tuples that force the cut vertex in $B_{i-1}\cap B_{i}$ to be the first vertex and the cut vertex in $B_{i-1}\cap B_{i}$ to be the last. In $\pi_{i}^{-}$ we do this the other way round. The partial order $\pi_{1}^{+}$ forces the cut vertex in $B_{1}$ to be the last vertex and $\pi_{1}^{-}$ forces it to be the first vertex. Similarly, $\pi_{k}^{+}$ forces the cut vertex of $B_{k}$ to be the first vertex and $\pi_{k}^{-}$ forces it to be the last vertex. Now we solve the MinPOHPP for all blocks $B_{i}$, once for $\pi_{i}^{+}$ and once for $\pi_{i}^{-}$. It is easy to see that the minimum cost ordered Hamiltonian path $\mathcal{P}$ of $G$ fulfilling the constraints of $\pi$ consists of minimum cost ordered Hamiltonian paths either for each instance $(B_{i},\pi_{i}^{+})$ or for each instance $(B_{i},\pi_{i}^{-})$. Thus, solving these instances is enough to solve the problem for the whole graph $G$. Note that the blocks of $G$ are also elements of $\mathcal{G}$ since $\mathcal{G}$ is hereditary.